# A Quantile Nelson-Siegel model

### v1

@@ Abstract
A widespread approach to modelling the interaction between macroeconomic variables and the yield curve relies on three latent factors usually interpreted as the level, slope, and curvature (Diebold et al., 2006). This approach is inherently focused on the conditional mean of the yields and postulates a dynamic linear model where the latent factors smoothly change over time.
However, periods of deep crisis, such as the Great Recession and the recent pandemic, have highlighted the importance of statistical models that account for asymmetric shocks and are able to forecast the tails of a variable’s distribution. A new version of the dynamic three-factor model is proposed to address this issue based on quantile regressions. The novel approach leverages the potential of quantile regression to model the entire (conditional) distribution of the yields instead of restricting to its mean.
An application to US data from the 1970s shows the significant heterogeneity of the interactions between financial and macroeconomic variables across different quantiles. Moreover, an out-of-sample forecasting exercise showcases the proposed method’s advantages in predicting the yield distribution tails compared to the standard conditional mean model. Finally, by inspecting the posterior distribution of the three factors during the recent major crises, new evidence is found that supports the greater and longer-lasting negative impact of the great recession on the yields compared to the COVID-19 pandemic.
Keywords: Bayesian inference; Nelson-Siegel model; Quantile regression; Time-varying parameters; Yield curve.
@@ 1 Introduction
Figure 1: Distributional term structure of the interest rate for June and December 2022.
@@
The main feature of the proposed TVP-QR-NS model is the coupling of quantile regression with time-varying three-factor Nelson-Siegel specification. In addition, to cope with the potentially high dimensionality of the latent dynamic factors, we use a precision sampler (Chan and Jeliazkov, 2009; Chan et al., 2020) that allows us to sample the entire path of the factors jointly without any loop. Overall, this combination provides a simple framework and computationally efficient method to investigate and forecast the term structure at every quantile level of interest.
The proposed method is then applied to investigate and forecast the US term structure of interest rates. Specifically, an in-sample analysis highlights the advantages of the quantile approach compared to the standard conditional mean Nelson-Siegel model. In particular, we are able to obtain the entire distribution of the three factors at different points in time, thus enabling a more detailed analysis of the effects of the recent crises (the Great Recession and the COVID-19 pandemic) on the term structure. Interestingly, these results, which could not be obtained using a conditional mean approach, reveal substantial differences between the impacts of the financial crisis and COVID-19. The former shifted leftwards the distribution of the short-term interest rate without meaningful impact on the long-term. Conversely, the pandemic had an initial significant adverse effect on both the short- and long-term rates, but the distribution in 2022 shifted rightwards compared to the pre-COVID one.
The advantages of the proposed TVP-QR-NS model are better highlighted in terms of tail risk yield forecasting. Compared to the conditional mean model, our approach provides better out-of-sample forecasts in both tails of the yield distribution, as measured by the quantile score. The difference is particularly striking on the left tail and consistent over time.
@@ 2.2 A Time-Varying Parameters Quantile NS Model
$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}\Big{)}+\beta_{3t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}-e^{-\lambda^{\tau}m}\Big{)},$ (3)
where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda^{\tau}\in\mathds{R}_{+}$ is a decay parameter. In the following, we shall assume $\lambda^{\tau}=\lambda$. Thus, by directly modelling the quantile function, we allow the yield curve factors to behave differently across the quantile levels.
Consider a series of observations at specific maturities $m\in\{1,\ldots,M\}$ and denote with $\mathbf{y}_{t}=(y_{t}(m_{1}),\ldots,y_{t}(m_{M}))^{\prime}$ the vector of yields at different maturity. Let $\mathbf{x}(m;\lambda)=\big{(}1,\frac{1-e^{-\lambda m}}{\lambda m},\frac{1-e^{-\lambda m}}{\lambda m}-e^{-\lambda m}\big{)}^{\prime}$, and $\mathbf{X}(\lambda)=\left[\mathbf{x}(m_{1};\lambda),\mathbf{x}(m_{2};\lambda),\ldots,\mathbf{x}(m_{M};\lambda)\right]^{\prime}\in\mathds{R}^{M\times 3}$. Then, the conditional quantile model can be written as
$Q^{\tau}_{t}=\mathbf{X}(\lambda)\bm{\beta}_{t}(\tau)$ (4)
@@
Differently from Coroneo et al. (2016) that proposed an approach related to ours in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
@@
$y_{t}(m)=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
@@
$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}(\tau)+w_{m,t}\sigma\theta_{\tau,1}+\sqrt{w_{m,t}}\sigma\theta_{\tau,2}u_{t}(m),\qquad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
where $w_{m,t}\sim\mathcal{E}xp(1)$ is exponential with rate one. To establish an equivalence between the mode of the asymmetric Laplace distribution and the optimisation of the standard quantile regression problem in Koenker and Bassett (1978), the parameters $\theta_{\tau,1},\theta_{\tau,2}$ are constrained to the following values (Kotz et al., 2001):
@@
Throughout the article, we follow the existing literature and fix the decay parameter at $\lambda=0.0609$ (e.g., see Diebold and Li, 2006). Moreover, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
@@
$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
where we recall the fixed values $\lambda=0.0609$ and $\sigma=1.0$. Denoting with $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{T})^{\prime}\in\mathds{R}^{T\times M}$ the matrix of observations, the joint density of the observables and auxiliary variables (conditional on the time-varying factors) is
$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{m,t}|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+w_{m,t}\sigma\theta_{\tau,1},\>w_{m,t}\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
@@ 3.1 Prior Specification
By leveraging the mixture representation of the asymmetric Laplace distribution, we obtain a conditionally linear Gaussian state space for the latent dynamic vector $\bm{\beta}_{t}$. Thus, we sample from the posterior full conditional distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
Let us define $e_{m,t}=y_{t}(m)-\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}$. Then, the posterior full conditional distribution of the auxiliary variable $w_{t}$ is obtained as
$w_{m,t}|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{m,t}^{-\frac{1}{2}}\exp\Big{\{}-\frac{1}{2}\Big{[}\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{m,t}+2w_{m,t}+w_{m,t}^{-1}\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\Big{]}\Big{\}},$ (12)
@@
$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
@@ 3.2 The precision sampler
$\mathbf{y}=\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
@@
$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+(\mathbf{H}S_{1})^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}(\mathbf{H}S_{1}),$
$\bm{\mu}_{b}$ $=K_{b}^{-1}\left(\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\right).$

### v2

@@ Abstract
We propose a novel framework for modeling the yield curve from a quantile perspective. Building on the dynamic Nelson–Siegel model of Diebold et al. (2006), we extend its traditional mean-based approach to a quantile regression setting, enabling the estimation of yield curve factors—level, slope, and curvature—at specific quantiles of the conditional distribution. A key advantage of our framework is its ability to characterize the entire conditional distribution of the yield curve across maturities and over time. In an empirical analysis of the U.S. term structure of interest rates, our method demonstrates superior out-of-sample forecasting performance, particularly in capturing the tails of the yield distribution—an aspect increasingly emphasized in the recent literature on distributional forecasting. In addition to its forecasting advantages, our approach reveals rich distributional features beyond the mean. In particular, we find that the dynamic changes in these distributional features differ markedly between the Great Recession and the COVID-19 pandemic period, highlighting a fundamental shift in how interest rate markets respond to distinct economic shocks.
Keywords: Nelson-Siegel; Yield Curve; Bayesian Markov Chain Monte-Carlo; Quantile regression.
@@ 1 Introduction
Figure 1: Distributional term structure of the interest rate for June (left) and December (right) 2022.
@@
The main feature of the proposed TVP-QR-NS model is the coupling of quantile regression with time-varying three-factor Nelson-Siegel specification. In addition, to cope with the potentially high dimensionality of the latent dynamic factors, we use a precision sampler (Chan and Jeliazkov, 2009; Chan et al., 2020) that allows us to sample the entire path of the factors jointly without any loop. Overall, this combination provides a simple framework and computationally efficient method to investigate and forecast the term structure at every quantile level of interest. Moreover, our novel econometric framework shares similarities with methodologies put forth by Korobilis and Schröder (2024, 2025). In the former, the authors propose a variational Bayes approach to estimate a quantile static factor model, while the latter extends the factor-augmented vector autoregression to a quantile framework. The primary distinction between our study and both of these lies in our specific focus on analysing the distribution of the yield curve. In contrast, their emphasis is on examining macroeconomic uncertainty within the US and macroeconomic risk across multiple countries.
The proposed method is applied to investigate and forecast the US term structure of interest rates, with particular emphasis on capturing the full distributional dynamics of key yield curve factors. To validate its utility, we conduct an out-of-sample forecasting exercise to evaluate the performance of the TVP-QR-NS model against standard conditional mean benchmarks widely used in the empirical literature. This study contributes to a growing body of research that focuses on forecasting macroeconomic variables from a distributional perspective (see Carriero et al. (2022); Clark et al. (2024); Lenza et al. (2025)). The results indicate that the TVP-QR-NS model consistently delivers superior tail risk forecasts relative to conventional benchmarks, particularly at longer forecast horizons and for long-term maturities. Moreover, employing the model confidence set methodology proposed by Hansen et al. (2011), we find that the TVP-QR-NS model is systematically retained within the confidence sets for longer maturities and long-term forecast horizons, while standard conditional mean models are frequently excluded, underscoring the robustness and reliability of our approach.
The in-sample analysis highlights the advantages of the quantile regression framework over the conventional conditional mean Nelson-Siegel model. Unlike mean-based frameworks, our method enables estimation of the entire distribution of the three Nelson-Siegel factors, as well as yields at selected maturities over time. This comprehensive characterization allows for a more detailed assessment of the effects of recent economic crises, specifically the Great Recession and the COVID-19 pandemic, on the term structure.
The results uncover critical insights that are not accessible through conditional mean models. Specifically, the impact of the Great Recession and COVID-19 on the distribution of short-term interest rates diverges substantially, with clear distributional shifts evident across the two crises. In contrast, the distribution of long-term rates during the Great Recession exhibited no significant shift in location, although an increase in left-skewness was observed.
Further analysis shows that, following the Great Recession, a pronounced location shift occurred across the short, medium, and long ends of the yield curve, with the 10-year maturity exhibiting increased left-skewness. Conversely, during the COVID-19 recession, the location shift was confined to the long end of the yield curve and was accompanied by heightened dispersion. Moreover, we find robust evidence of quantile-dependent dynamics in the relationship between yield curve factors and macroeconomic conditions, underscoring the importance of capturing heterogeneity across the distribution.
@@ 2.2 A Time-Varying Parameters Quantile NS Model
$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\left(\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}\right)+\beta_{3t}(\tau)\left(\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}-e^{-\lambda(\tau)m}\right),$ (3)
where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda(\tau)\in\mathds{R}_{+}$ is a decay parameter. Thus, by directly modelling the quantile function, we allow the yield curve factors to behave differently across the quantile levels.
Consider a series of observations at specific maturities $m\in\{1,\ldots,M\}$ and denote with $\mathbf{y}_{t}=(y_{t}(m_{1}),\ldots,y_{t}(m_{M}))^{\prime}$ the vector of yields at different maturity. Let $\mathbf{x}(m;\lambda(\tau))=\big{(}1,\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m},\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}-e^{-\lambda(\tau)m}\big{)}^{\prime}$, and $,\mathbf{X}(\lambda(\tau))=[\mathbf{x}(m_{1};\lambda(\tau)),\mathbf{x}(m_{2};\lambda(\tau)),$$\ldots,\mathbf{x}(m_{M};\lambda(\tau))]^{\prime}\in\mathds{R}^{M\times 3}$. Then, the conditional quantile model can be written as
$Q^{\tau}_{t}=\mathbf{X}(\lambda(\tau))\bm{\beta}_{t}(\tau)$ (4)
@@
Differently from Coroneo et al. (2016) that proposed an approach related to our in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
@@
$y_{t}(m)=\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
@@
$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+w_{t}(m)\sigma\theta_{\tau,1}+\sqrt{w_{t}(m)}\sigma\theta_{\tau,2}u_{t}(m),\quad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
where $w_{t}(m)\sim\mathcal{E}xp(1)$ is exponential with rate one. To establish an equivalence between the mode of the asymmetric Laplace distribution and the optimisation of the standard quantile regression problem in Koenker and Bassett (1978), the parameters $\theta_{\tau,1},\theta_{\tau,2}$ are constrained to the following values (Kotz et al., 2001):
@@
In the existing literature, the decay parameter is often fixed at $\lambda=0.0609$ (e.g., Diebold and Li, 2006). In contrast, we adopt an agnostic approach by estimating $\lambda(\tau)$ separately for each quantile, and assume a uniform prior distribution, $\lambda(\tau)\sim\mathcal{U}(0,0.3)$ (see the appendix for further details). Given that the conditional posterior distribution of $\lambda(\tau)$ is non-standard, we employ a Griddy-Gibbs sampling step to draw from its posterior. Finally, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
@@
$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
Denoting with $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{T})^{\prime}\in\mathds{R}^{T\times M}$ the matrix of observations, the joint density of the observables and auxiliary variables (conditional on the time-varying factors) is
$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{t}(m)|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+w_{t}(m)\sigma\theta_{\tau,1},\>w_{t}(m)\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
@@ 3.1 Prior Specification
By leveraging the mixture representation of the asymmetric Laplace distribution, we obtain a conditionally linear Gaussian state space for the latent dynamic vector $\bm{\beta}_{t}$. Thus, we sample the full conditional posterior distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
Let us define $e_{t}(m)=y_{t}(m)-\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}$. Then, the full conditional posterior distribution of the auxiliary variable $w_{t}(m)$ is obtained as
$w_{t}(m)|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{t}(m)^{-\frac{1}{2}}\exp\left\{-\frac{1}{2}\left[\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{t}(m)+2w_{t}(m)+w_{t}(m)^{-1}\frac{e_{t}(m)^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\right]\right\},$ (12)
@@
$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{t}(m)^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
@@ 3.2 The precision sampler
$\mathbf{y}=\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\right)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
@@
$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\big{(}\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\big{)}^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+S_{1}\mathbf{H}^{\prime}\big{(}\mathbf{I}_{T}\otimes\Omega^{-1}\big{)}\mathbf{H}S_{1},$
$\bm{\mu}_{b}$ $=K_{b}^{-1}\Big{(}\big{(}\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\big{)}^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\big{(}\mathbf{I}_{T}\otimes\Omega^{-1}\big{)}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\Big{)}.$
3.3 Monotonicity
Traditional estimation of quantile regression, including our strategy, involves fitting the quantile NS model separately at each quantile level, treating each quantile as an independent modeling task. Although this approach provides flexibility and ease of implementation, it overlooks the intrinsic monotonicity that should characterize conditional quantile functions by construction. As a result, the estimated quantile curves may violate the logical ordering across quantiles, potentially leading to internal inconsistencies and undermining the interpretability of the distributional forecasts.
To ensure monotonicity across quantile levels $\tau\in\left\{\tau_{1},...,\tau_{K}\right\}$, it is essential to move beyond separate estimations and instead estimate the quantile NS models jointly across $\tau$. This joint estimation framework requires imposing explicit monotonicity constraints across quantiles,
$\mathbf{x}\big{(}m;\lambda(\tau_{k})\big{)}^{\prime}\bm{\beta}_{t}(\tau_{k})\leq\mathbf{x}\big{(}m;\lambda(\tau_{k+1})\big{)}^{\prime}\bm{\beta}_{t}(\tau_{k+1})$
for $m=1,\ldots,M,k=1,\ldots,K$ and $t=1,\ldots T$, thus preserving the fundamental ordering structure inherent to conditional quantile functions. However, this approach introduces a large number of cross-quantile constraints, and the resulting optimization problem becomes highly complex and computationally burdensome, to the point where joint estimation under monotonicity constraints can be practically infeasible in its full form. This poses a major challenge for implementing a theoretically coherent yet computationally tractable distributional forecasting framework. To construct a coherent predictive density from the individually estimated quantile NS models, we adopt the re-ordering technique proposed by Chernozhukov et al. (2010), a standard approach in the literature for addressing the issue of quantile crossing.
In the context of the NS model, the yield curve is endowed with additional structural interpretation that can inform both estimation and inference. Specifically, the level factor in the NS model captures the long-run component of interest rates and asymptotically corresponds to the yield at extremely long maturities—essentially representing the ultimate long-term rate. Conversely, the sum of the level and slope factors determines the instantaneous (short-term) interest rate, reflecting the yield at the shortest maturity. Based on these insights, we can potentially impose the following monotonicity constraints
$\bm{\beta}_{t,1}(\tau_{k})\leq\bm{\beta}_{t,1}(\tau_{k+1})$ (17)
and
$\bm{\beta}_{t,1}(\tau_{k})+\bm{\beta}_{t,2}(\tau_{k})\leq\bm{\beta}_{t,1}(\tau_{k+1})+\bm{\beta}_{t,2}(\tau_{k+1}).$ (18)
Instead of directly imposing these monotonicity constraints within a joint estimation framework—such as through constrained MCMC, which would be computationally intensive—we adopt a more practical post-processing strategy. Specifically, we formulate a quadratic programming problem that adjusts the unconstrained estimates of $\bm{\beta}_{t,1}(\tau_{k})$ and $\bm{\beta}_{t,2}(\tau_{k})$ across quantile levels. The objective is to find new values that remain as close as possible to the originally estimated posterior mean, $\widehat{\bm{\beta}}_{t,1}$ and $\widehat{\bm{\beta}}_{t,2}$, while satisfying the monotonicity constraints. This post-estimation correction ensures internal consistency and preserves the structural interpretation of the NS model, while avoiding the computational burden of imposing constraints during the sampling phase.